OpenStudy (anonymous):

If M(x)=2f(x)/g(x), what is the value of M'(-1)? We have to use the quotient rule, and there's a table. I'll try to write what's needed to do the problem. f(x) when x=-1 is 3, f'(x) when x=-1 is -2, g(x) when x=-1 is 1, and g'(x) when x=-1 is also 1. Please show steps! I got M'(-1)=-10, but I'm unsure. Thanks!

myininaya (myininaya):

You need to apply quotient rule for the f/g part.

myininaya (myininaya):

recall \[(\frac{f(x)}{g(x)})'=\frac{f'(x) \cdot g(x)-f(x) \cdot g'(x)}{(g(x))^2}\]

myininaya (myininaya):

now just multiply 2 on both sides to get (2f/g)'

myininaya (myininaya):

and since you are asked to find M'(-1) replace the x's with -1

myininaya (myininaya):

then use f(-1)=3 f'(-1)=-2 g(-1)=1 and g'(-1)=1

myininaya (myininaya):

then use order of operations to simiplify